# One Backpropagation in Two Tower Recommendation Models

### v3

@@ Abstract.
isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommender systems have been developing for decades to help people dealing with the issue of information overload by recommending a user only a few of his mostly interested items [3; 2]. With the increasing concern on privacy, users and items are only equipped with unique identifiers without more descriptions, and various users behaviors, such as view, click, purchase and etc., are simplified as binary interactions between users and items, i.e., interacted or uninteracted. Recommendation only based on such binary interactions is also well known as the One Class Collaborative Filtering (OCCF) [30; 15]: How to recommend a sorted list consisting of uninteracted items for each user based only on a user-item interaction matrix?
Extensive research efforts have been devoted to dealing with the OCCF problem [14; 38; 25; 43; 23; 6; 21; 5; 4; 19; 22; 42]. The key to success lies in how to learn high quality users’ and items’ representations from their unique identifiers and binary interactions. A user-item predicted rating matrix can then be computed based on users’ and items’ representations to produce recommendation lists. Most proposed algorithms for the OCCF problem have adopted a kind of two tower recommendation model [26; 16; 44; 36], in which one tower for encoding users’ representations and another for encoding items’ representations. Furthermore, these algorithms have employed the contrastive learning technique [7; 11; 17; 35] to differentiate interacted and uninteracted items in encoding representations.
We can identify four building modules in most two tower recommendation models, namely, user-item encoding, negative sampling, loss computing and backpropagation updating. Many recommendation algorithms have been proposed with the focus of improving the first three modules. For user-item encoding, many advanced yet sophisticated neural networks have been designed with the aim of better encoding collaborative information from the interaction matrix [13; 40; 43; 23; 6; 21]. For negative sampling, the main objective is to select those true negatives from uninteracted, i.e., an uninteracted item truly disliked by a user. For loss computing, the InfoNCE [29] has become a rather popular contrastive learning loss function, and some algorithms have proposed to further improve it, say for example, by using different similarity computation functions [14; 18]
@@
In this paper, we challenge such an equal treating assumption in two backpropagation for model training. We first argue that even though items are equipped with only identifiers, they are indeed also with some intrinsic attributes (like a movie genre and director), even not exposed in the OCCF scenarios. We next argue that a collaborative filtering model could be capable of learning such unexposed attributes as encoded in the learned items’ representations, and as such, items can be clustered into latent types. We also argue that a user’s interests could be diverse, embodied in that his interacted items normally are distributed in different latent types. We note that in the batch-based training phase, the gradient direction would be dominated by those items with a same type. As such, using the two backpropagation may lead to the learned representation of a user too much favoring those of the same type items, while neglecting items of other types.
@@
Motivated from such considerations, we propose one backpropagation (OneBP) in two tower recommendation models. The core idea is rather simple: We keep the normal gradient backpropagation for the item encoding tower, but cut off the backpropagation for the user encoding tower. Instead, we propose a moving-aggregation strategy to update a user encoding in each training epoch. In particular, the user encoding is updated by the aggregation over his previous one and his just-updated interacted item’s encoding. Except the new backpropagation strategy, we implement other three modules with the most straightforward choices, namely, the projection matrix-based user-item encoding, random negative sampling and plain InfoNCE loss. We experiment our OneBP model on four widely used public datasets. Results validate its superiority over representative state-of-the-art models in terms of large improvement margins on recommendation performance, as well as its computation efficiency.
@@ 2. Two Tower Recommendation Models
User-Item Encoding: This module consists a user encoder to encode a user from its one-hot $M$-dimensional identifier $u_{i}$ to a real-valued $d$-dimensional ($d\ll M$) representation $\mathbf{u}_{i}$, and an item encoder to encode an item $v_{j}$ to a $d$-dimensional representation $\mathbf{v}_{j}$:
@@
Negative Sampling: Due to the nature of binary interactions, the contrastive learning has been serving as the main solution approach for the OCCF problem. In the OCCF setting, a positive sample is from interactions $(u_{i},v_{i}^{+})$ between a user $u_{i}$ and one of his interacted items; While a negative sample $(u_{i},v_{i}^{-})$ is from one of his uninteracted items. In recommendation systems, uninteracted items are much more than interacted ones, and negative sampling is needed to select a number of uninteracted items (say $N_{s}\ll N$):
@@
Notice that an uninteracted item could be the one not seen before by a user but not an uninterested item to this user. As such, many negative sampling algorithms have been proposed to select high-quality uninteracted items as negatives, such as the PNS [27; 8; 37], DNS [46] and etc.
@@
The commonly used similarity function includes cosine, inner product function. Let $r_{ij}^{+}$ ($r_{ij}^{-}$) denote the rating between a user $u_{i}$ and his positive item (negative item). In the OCCF problem, the popular contrastive loss function is the following InforNCE loss [29]:
@@
Backpropagation Updating: During the model training, the mostly used technique is the gradient backpropagation, which applies the gradients calculated from the loss to update the inputs of the encoders. To the best of our knowledge, all existing two tower models use the same updating strategy to update both the user encoder and item encoder. Although different encoders have different structures and parameters, the objective of gradient backpropagation is to update the user representation and item representation after encoders in each iteration. As such, we can summarize the backpropagation updating in one training epoch as follows:
@@
where $\alpha$ is the learning rate.
@@ 3.1. Encoder-based method
Recently, many advanced neural networks have been designed as encoders [13; 40; 43; 23; 6; 21], among which the NGCF [40] and LightGCN [43] are two representative ones of discriminative model. The NGCF [40] applies a graph neural network to recursively propagates and aggregates collaborative signals in the user-item interaction graph in order to capture high-order structural information. The LightGCN [43] argues that the feature transformation and non-linear activation in NGCF do not positively impact collaborative filtering. As such, the LightGCN significantly simplifies the graph neural network by only retaining the most crucial components, leading to improved performance.
@@ 3.2. Negative Sampling method
For negative sampling, the challenge is from the nature of unlabeled data in the OCCF problem. Many approaches have focused on selecting high-quality true negatives [32; 46; 14; 38; 25], yet some focus on other criteria in negative sampling, such as mitigating popularity bias [10; 27; 24]. In negative sampling, the probability of selecting an uninteracted item can be predefined (static) or adjusted in each iteration (dynamic).
The RNS (random negative sampling) [32; 43; 45] is the widely used static negative sampling, which selects uninteracted items with equal probabilities. The PNS (popularity-based negative sampling) [27; 8; 37] computes the sampling probability for an uninteracted item to a user proportional to the item popularity that is defined as the interaction frequency of this item across all users in the training dataset.
@@ 3.3. Loss Function
Although the InfoNCE in Eq. (2) has been becoming a kind of standard loss function for contrastive learning, some have proposed to further improve it from two viewpoints: One is to modify the similarity calculation function [14; 5; 4; 19; 22; 42], and another is to alleviate the adverse impact of false negatives) [11; 33].
@@ 4.1. The common modules
Projection-based User-Item Encoding: We use a very simple projection for encoding users and items. Let $\mathbf{W}_{u}\in\mathbb{R}^{M\times k}$ denote the user projection matrix which converts a user one-hot identifier vector into a $d$-dimensional real-valued representation. Simply put, the $i$-th row of $\mathbf{W}_{u}$ is the user $u_{s}$’s representation. Similarly, $\mathbf{W}_{v}\in\mathbb{R}^{N\times k}$ is the item projection matrix. The two projection matrices $\mathbf{W}_{u}$ and $\mathbf{W}_{v}$ are subject to be updated in the model training phase. Before model training, they are randomly initialized.
Random Negative Sampling: We adopt the most straightforward random negative smpling: For each positive sample $(u_{i},v_{i}^{+})$, we randomly select his $N_{s}$ uninteracted items as negative samples $(u_{i},v_{i}^{-})$. We notice that although many advanced negative sampling techniques can be used to promote contrastive learning, the choice of the random one is with a significant merit of fast computation.
@@ 4.2. One Backpropagation Updating
In this paper, we propose to use loss gradient backpropgation for updating items’ representation only, and propose to update users’ representations by our moving-aggregation strategy. We call our updating strategy as ”OneBP” for brevity. Fig. 2 compares our OneBP with the two backpropagation. The core idea is to cut off the gradient flow towards users’ encodings.
We detail the execution of our OneBP strategy for one user in one training epoch as follows: Let $(u_{i},v_{j},\{v_{k}\})$ denote the selected user $u_{i}$ with his interacted item $v_{j}$ as positive and his uninteracted items $\{v_{k}\}$ as negatives. Our OneBP updates $\mathbf{u}_{i}$, $\mathbf{v}_{j}$, and $\mathbf{v}_{k}$ as follows:
@@
Compared with the two backpropagation, our OneBP does not use gradient backpropagation for updating users’ representations. Instead, Eq. (6) indicates that the user encoding is updated by the aggregation over his previous one and his just-updated interacted item’s encoding in a training epoch.
@@ 4.3.2. Decoupling the Impact of False Negative Samples
We note that although contrastive learning is a powerful technique in many applications, it faces the challenge of false negative sample in the OCCF problem. That is, an unlabeled item is a user interested one, but it is not interacted simply due to not yet exposed to this user. If such an unlabeled item is selected as a negative sample for contrastive loss computation, it can lead to update user encoding in an incorrect direction. This issue has been extensively studied in the literature [34; 20; 9; 12]. Although many algorithms have been proposed to alleviate the adverse impact of false negative samples [11; 33; 25; 41; 31; 9], the existing two backpropagation cannot fully address this issue as its loss gradient is still backpropagated to update user encodings. In contrast, our OneBP only applies his interacted items’ encodings to update a user encoding, which can effectively decouple the impact of false negative samples.
@@ 5.1. Experiment Settings
Dataset: We conduct experiments on four widely used public datasets, including the MovieLens-100k, MovieLens-1M  [^1] MovieLens: https://grouplens.org/datasets/movielens/, Gowalla  [^2] Gowalla: https://paperswithcode.com/dataset/gowalla and Yelp2018  [^3] Yelp2018: https://www.yelp.com/dataset. Table 1 presents the statistics of the four datasets. We note that the Gowalla and Yelp2018 are often regarded as large datasets with serious interaction sparsity.
@@
Competitors: We compare our proposed OneBP with the following three groups of competitors, as introduced in Section 3. (1) The first encoder group focuses on the design of user and item encoder, including the MF [18], NGCF [40], and LightGCN[43]. (2) The second negative group focuses on negative sampling yet with the simple MF encoder, including the PNS [27; 8; 37], DNS [46], HarS [14], and 2stS [38]. (3) The third loss group focuses on the loss function, including the CML [14], SFCML [4], DPCML [5], DCL [11], HCL [33], and BUIR [19].
[…]
Parameter Settings: To make fair comparisons, we use the same number of negative samples for all algorithms. More detailed hyperparameter settings are documented in the README file of the code. The codes is implemented with Pytorch  [^4] Codes are released at: https://anonymous.4open.science/r/OneBP.
@@ 5.4. Hyper-parameter Study
Our OneBP contains only one hyperparameter $\beta$ to balance the weight between a user encoding in last iteration and his interacted item encoding updated in this iteration, c.f., Eq. (6). Note that the number of negatives $N_{s}$ is a common hyperparameter in contrastive learning. We conduct experiments to analyze the effects of the two hyperparameters.
@@ 5.5. Exploration Study
Table 4. Statistics of genre on Movielens_100k.
Category unknown Action Adventure Animation Children’s Comedy Crime Documentary Drama Fantasy
Amount 2 251 135 42 122 505 109 50 725 22
Category Film-Noir Horror Musical Mystery Romance Sci-Fi Thriller War Western Total
Amount 24 92 56 61 247 101 251 71 27 2893
Table 5. Capability of collaborative filtering models in identifying latent types.
Top5 classification performance Top10 classification performance
Category Precision Recall F1-score Category Precision Recall F1-score
TwoBP OneBP TwoBP OneBP TwoBP OneBP TwoBP OneBP TwoBP OneBP TwoBP OneBP
unknown 0 0 0 0 0 0 unknown 0.1 0 0.5 0 0.167 0
Action 1 1 0.02 0.02 0.039 0.039 Action 1 1 0.04 0.04 0.077 0.077
Adventure 0.8 0.8 0.03 0.03 0.057 0.057 Adventure 0.6 0.7 0.044 0.052 0.083 0.097
Animation 0.8 1 0.095 0.119 0.17 0.213 Animation 0.7 0.9 0.167 0.214 0.269 0.346
Children’s 1 1 0.041 0.041 0.079 0.079 Children’s 1 1 0.082 0.082 0.152 0.152
Comedy 0.8 1 0.008 0.01 0.016 0.02 Comedy 0.7 1 0.014 0.02 0.027 0.039
Crime 0.4 0.6 0.018 0.028 0.035 0.053 Crime 0.2 0.6 0.018 0.055 0.034 0.101
Documentary 0.2 0.2 0.02 0.02 0.036 0.036 Documentary 0.1 0.1 0.02 0.02 0.033 0.033
Drama 0.8 1 0.006 0.007 0.011 0.014 Drama 0.6 1 0.008 0.014 0.016 0.027
Fantasy 0.4 0.4 0.091 0.091 0.148 0.148 Fantasy 0.2 0.4 0.091 0.182 0.125 0.25
Film-Noir 0.4 0.6 0.083 0.125 0.138 0.207 Film-Noir 0.3 0.5 0.125 0.208 0.176 0.294
Horror 1 1 0.054 0.054 0.103 0.103 Horror 1 1 0.109 0.109 0.196 0.196
Musical 0.6 1 0.054 0.089 0.098 0.164 Musical 0.5 1 0.089 0.179 0.152 0.303
Mystery 0 0.2 0 0.016 0 0.03 Mystery 0.1 0.4 0.016 0.066 0.028 0.113
Romance 0.4 0.8 0.008 0.016 0.016 0.032 Romance 0.4 0.9 0.016 0.036 0.031 0.07
Sci-Fi 0.6 0.8 0.03 0.04 0.057 0.075 Sci-Fi 0.6 0.8 0.059 0.079 0.108 0.144
Thriller 0.8 0.8 0.016 0.016 0.031 0.031 Thriller 0.6 0.7 0.024 0.028 0.046 0.054
War 0.6 0.8 0.042 0.056 0.079 0.105 War 0.4 0.6 0.056 0.085 0.099 0.148
Western 0.2 0 0.037 0 0.062 0 Western 0.1 0.2 0.037 0.074 0.054 0.108
Overall 0.568 0.684 0.019 0.022 0.036 0.044 Overall 0.484 0.674 0.032 0.044 0.06 0.083
Improve 20.42$\uparrow$ 15.79$\uparrow$ 22.22$\uparrow$ 39.26$\uparrow$ 37.50$\uparrow$ 38.33$\uparrow$
Top20 classification performance Top50 classification performance
Category Precision Recall F1-score Category Precision Recall F1-score
TwoBP OneBP TwoBP OneBP TwoBP OneBP TwoBP OneBP TwoBP OneBP TwoBP OneBP
unknown 0.05 0 0.5 0 0.091 0 unknown 0.02 0.02 0.5 0.5 0.038 0.038
Action 1 1 0.08 0.08 0.148 0.148 Action 0.86 1 0.171 0.199 0.286 0.332
Adventure 0.6 0.7 0.089 0.104 0.155 0.181 Adventure 0.48 0.64 0.178 0.237 0.259 0.346
Animation 0.65 0.85 0.31 0.405 0.419 0.548 Animation 0.4 0.48 0.476 0.571 0.435 0.522
Children’s 0.85 1 0.139 0.164 0.239 0.282 Children’s 0.76 0.86 0.311 0.352 0.442 0.5
Comedy 0.6 1 0.024 0.04 0.046 0.076 Comedy 0.42 0.96 0.042 0.095 0.076 0.173
Crime 0.15 0.55 0.028 0.101 0.047 0.171 Crime 0.12 0.46 0.055 0.211 0.075 0.289
Documentary 0.1 0.1 0.04 0.04 0.057 0.057 Documentary 0.08 0.04 0.08 0.04 0.08 0.04
Drama 0.75 0.75 0.021 0.021 0.04 0.04 Drama 0.72 0.74 0.05 0.051 0.093 0.095
Fantasy 0.15 0.35 0.136 0.318 0.143 0.333 Fantasy 0.12 0.2 0.273 0.455 0.167 0.278
Film-Noir 0.3 0.35 0.25 0.292 0.273 0.318 Film-Noir 0.14 0.2 0.292 0.417 0.189 0.27
Horror 1 1 0.217 0.217 0.357 0.357 Horror 0.82 0.9 0.446 0.489 0.577 0.634
Musical 0.35 0.5 0.125 0.179 0.184 0.263 Musical 0.28 0.3 0.25 0.268 0.264 0.283
Mystery 0.1 0.3 0.033 0.098 0.049 0.148 Mystery 0.06 0.3 0.049 0.246 0.054 0.27
Romance 0.3 0.9 0.024 0.073 0.045 0.135 Romance 0.18 0.7 0.036 0.142 0.061 0.236
Sci-Fi 0.6 0.8 0.119 0.158 0.198 0.264 Sci-Fi 0.36 0.6 0.178 0.297 0.238 0.397
Thriller 0.65 0.7 0.052 0.056 0.096 0.103 Thriller 0.48 0.72 0.096 0.143 0.159 0.239
War 0.4 0.4 0.113 0.113 0.176 0.176 War 0.26 0.36 0.183 0.254 0.215 0.298
Western 0.15 0.15 0.111 0.111 0.128 0.128 Western 0.1 0.12 0.185 0.222 0.13 0.156
Overall 0.461 0.6 0.06 0.079 0.107 0.139 Overall 0.351 0.505 0.115 0.166 0.173 0.25
Improve 30.15$\uparrow$ 31.67$\uparrow$ 29.91$\uparrow$ 43.87$\uparrow$ 44.35$\uparrow$ 44.51$\uparrow$
  • 1
  • 2
    The colors represent the classification performance of one model over the other, where orange indicates the phenomenon that OneBP’s classification performance is better than TwoBP’s, and green represents the opposite situation.
Figure 6. The experimental process of model classification performance exploration based on Movielens_100k dataset
@@
We next explore our arguments on items’ encodings and latent types. An interesting question is about whether the learned items’ encodings are with the capability of differentiating latent types. To this end, we first resort to the ’genre’ attribute in the Movielens100k dataset as the ground truth labels for items’ types and there are 19 genre types, Table 4 shows its statistics infomation. For each genre, we calculate its encoding as the averaged encodings of those items with the same genre label, as shown in the upper part of Fig. 6. We then examine whether the top-$k$ items closest to each genre in the encoding space are with the same genre label, as shown in the lower part of Fig. 6. Precision, Recall, and F1 scores are employed as classification metrics, where high Precision indicates a good capability of identifying item types. Table 5 presents the experimental results. In most genres, OneBP exhibits stronger classification capability compared to TwoBP. Particularly, in genres with a large number of items, OneBP significantly outperforms TwoBP in terms of classification performance. From the overall average classification performance perspective, OneBP also outperforms TwoBP by a large margin.
However, data often lack labels in the OCCF problem, making it impossible to measure the quality of learned encodings as well as the diversity of user preferences. For items without labels, we resort to the KMeans clustering algorithm [28] to cluster items’ encodings into six clusters, where one item belongs to one cluster. We use the t-SNE dimensionality reudction [39] to map items’ encodings into a two-dimensional plane, where colors indicate cluster indexes, i.e., latent types.
Fig. 4 plots the distribution of projected items’ encodings for both the training and testing dataset, where a user encoding is also projected and his interacted items in the training dataset are marked as rectangles and with dark green color. Fig. 4 (a) presents the distribution of items’ encodings at the random initialization phase. We can observe that items with different latent types are almost mixed and there are no clear cluster divisions. Fig. 4 (b)(c)(d) present the distribution of items’ encodings after the training of a two tower model. It can be clearly seen that items of the same color (indicating of the same latent type) are closer to each others as a cluster, and the divisions of different clusters are clearly observable. This validates our arguments to some extent, that is, a two tower recommendation model is capable of learning items’ intrinsic yet unexposed attributes as encoded in the learned items’ encodings, and items can be clustered into laten types with observable cluster divisions.
Fig. 4 (b) displays the clustering results by using one backpropagation only for the user encoding update. It can be observed that the learned items’ encodings do not present a kind of uniform dispersion in the space, and clusters’ divisions are also not that clear, as compared with Fig. 4 (c) and (d). Furthermore, the interacted items by a user in the training dataset are too much concentrated in one cluster, and the user encoding is farther away from his interacted items, also compared with Fig. 4 (c) and (d). These observations indicate that this user backpropagation training strategy may not be capable of learning high quality representations for items and for users, so its recommendation performance suffers.

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recommender systems have been developed for decades to help people dealing with the issue of information overload by recommending a user only a few of his/her mostly interested items [1; 2]. With the increasing concern on privacy, users and items are only equipped with unique identifiers without more descriptions, and various users behaviors, such as view, click, purchase and etc., are simplified as binary interactions between users and items, i.e., interacted or uninteracted. Recommendation only based on such binary interactions is also well known as the One Class Collaborative Filtering (OCCF) [3; 4]: How to recommend a sorted list consisting of uninteracted items for each user based only on a user-item interaction matrix?
Extensive research efforts have been devoted to dealing with the OCCF problem [5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15; 16]. The key to success lies in how to learn high quality users’ and items’ representations from their unique identifiers and binary interactions. A user-item predicted rating matrix can then be computed based on users’ and items’ representations to produce recommendation lists. Most proposed algorithms for the OCCF problem have adopted a kind of two tower recommendation model [17; 18; 19; 20], in which one tower for encoding users’ representations and another for encoding items’ representations. Furthermore, these algorithms have employed the contrastive learning technique [21; 22; 23; 24] to differentiate interacted and uninteracted items in encoding representations.
We can identify four building modules in most two tower recommendation models, namely, user-item encoding, negative sampling, loss computing and backpropagation updating. Many recommendation algorithms have been proposed with the focus of improving the first three modules. For user-item encoding, many advanced yet sophisticated neural networks have been designed with the aim of better encoding collaborative information from the interaction matrix [25; 26; 8; 9; 10; 11]. For negative sampling [27; 28; 29; 5; 6; 7; 30], the main objective is to select those true negatives from uninteracted, i.e., an uninteracted item truly disliked by a user. For loss computing, the InfoNCE [31] has become a rather popular contrastive learning loss function, and some algorithms have proposed to further improve it, say for example, by using different similarity computation functions [5; 32].
@@
In this paper, we challenge such an equal treating assumption in two backpropagation for model training. We first argue that even though items are equipped with only identifiers, they are indeed also with some intrinsic attributes (like a movie genre and director), even not exposed in the OCCF scenarios. We next argue that a collaborative filtering model could be capable of learning such unexposed attributes as encoded in the learned items’ representations, and as such, items can be clustered into latent types. We also argue that a user’s interests could be diverse, embodied in that his/her interacted items normally are distributed in different latent types. We note that in the batch-based training phase, the gradient direction would be dominated by those items with a same type. As such, using the two backpropagation may lead to the learned representation of a user too much favoring those of the same type items, while neglecting items of other types.
@@
Motivated from such considerations, we propose one backpropagation (OneBP) in two tower recommendation models. The core idea is rather simple: We keep the normal gradient backpropagation for the item encoding tower, but cut off the backpropagation for the user encoding tower. Instead, we propose a moving-aggregation strategy to update a user encoding in each training epoch. In particular, the user encoding is updated by the aggregation over his/her previous encoding and his/her just-interacted item’s encoding. Except the new backpropagation strategy, we implement other three modules with the most straightforward choices, namely, the projection matrix-based user-item encoding, random negative sampling and plain InfoNCE loss. We experiment our OneBP model on four widely used public datasets. Results validate its superiority over representative state-of-the-art models in terms of large improvement margins on recommendation performance, as well as its computation efficiency.
@@ 2. Two Tower Recommendation Models
User-Item Encoding: This module consists of a user encoder to encode a user from its one-hot $M$-dimensional identifier $u_{i}$ to a real-valued $d$-dimensional ($d\ll M$) representation $\mathbf{u}_{i}$, and an item encoder to encode an item $v_{j}$ to a $d$-dimensional representation $\mathbf{v}_{j}$:
@@
Negative Sampling: Due to the nature of binary interactions, the contrastive learning has been serving as the main solution approach for the OCCF problem. In the OCCF setting, a positive sample is from interactions $(u_{i},v_{i}^{+})$ between a user $u_{i}$ and one of his/her interacted items; While a negative sample $(u_{i},v_{i}^{-})$ is from one of his/her uninteracted items. In recommendation systems, uninteracted items are much more than interacted ones, and negative sampling is needed to select a number of uninteracted items (say $N_{s}\ll N$):
@@
Notice that an uninteracted item could be the one not seen before by a user but not an uninterested item to this user. As such, many negative sampling algorithms have been proposed to select high-quality uninteracted items as negatives, such as the PNS [33; 27; 34], DNS [29] and etc.
@@
The commonly used similarity function includes cosine, inner product function. Let $r_{ij}^{+}$ ($r_{ik}^{-}$) denote the rating between a user $u_{i}$ and his/her positive item $v_{j}$ (negative item $v_{k}$ sampled from his/her uninteracted item sets $\mathcal{V}_{u_{i}}^{-}$). In the OCCF problem, the popular contrastive loss function is the following InfoNCE loss [31]:
@@
Backpropagation Updating: During the model training, the most used technique is the gradient backpropagation, which applies the gradients calculated from the loss to update the inputs of the encoders. To the best of our knowledge, all existing two tower models use the same updating strategy to update both the user encoder and item encoder. Although different encoders have different structures and parameters, the objective of gradient backpropagation is to update the user representation and item representation after encoders in each iteration. As such, we can summarize the backpropagation updating in one training epoch as follows:
@@
where $\alpha$ is the learning rate and $\nabla{\mathcal{L}}$ is a partial derivative w.r.t. the parameters being updated.
@@ 3.1. Encoder-based method
Recently, many advanced neural networks have been designed as encoders [25; 26; 8; 9; 10; 11], among which the NGCF [26] and LightGCN [8] are two representative ones of discriminative model. The NGCF [26] applies a graph neural network to recursively propagates and aggregates collaborative signals in the user-item interaction graph in order to capture high-order structural information. The LightGCN [8] argues that the feature transformation and non-linear activation in NGCF do not positively impact collaborative filtering. As such, the LightGCN significantly simplifies the graph neural network by only retaining the most crucial components, leading to improved performance.
@@ 3.2. Negative Sampling method
For negative sampling, the challenge is from the nature of unlabeled data in the OCCF problem. Many approaches have focused on selecting high-quality true negatives [28; 29; 5; 6; 7], yet some focus on other criteria in negative sampling, such as mitigating popularity bias [30; 33; 35]. In negative sampling, the probability of selecting an uninteracted item can be predefined (static) or adjusted in each iteration (dynamic).
The RNS (random negative sampling) [28; 8; 36] is the widely used static negative sampling, which selects uninteracted items with equal probabilities. The PNS (popularity-based negative sampling) [33; 27; 34] computes the sampling probability for an uninteracted item to a user proportional to the item popularity that is defined as the interaction frequency of this item across all users in the training dataset.
@@ 3.3. Loss Function
Although the InfoNCE in Eq. (2) has been becoming a kind of standard loss function for contrastive learning, some have proposed to further improve it from two viewpoints: One is to modify the similarity calculation function [5; 12; 13; 14; 15; 16], and another is to alleviate the adverse impact of false negatives) [22; 37].
@@ 4.1. The common modules
Projection-based User-Item Encoding: We use a very simple projection for encoding users and items. Let $\mathbf{W}_{u}\in\mathbb{R}^{M\times k}$ denote the user projection matrix which converts a user one-hot identifier vector into a $d$-dimensional real-valued representation. Simply put, the $i$-th row of $\mathbf{W}_{u}$ is the user $u_{i}$’s representation. Similarly, $\mathbf{W}_{v}\in\mathbb{R}^{N\times k}$ is the item projection matrix. The two projection matrices $\mathbf{W}_{u}$ and $\mathbf{W}_{v}$ are subject to be updated in the model training phase. Before model training, they are randomly initialized.
Random Negative Sampling: We adopt the most straightforward random negative sampling: For each positive sample $(u_{i},v_{i}^{+})$, we randomly select his/her $N_{s}$ uninteracted items as negative samples $(u_{i},v_{i}^{-})$. We notice that although many advanced negative sampling techniques can be used to promote contrastive learning, the choice of the random one is with a significant merit of fast computation.
@@ 4.2. One Backpropagation Updating
In this paper, we propose to use loss gradient backpropagation for updating items’ representation only, and propose to update users’ representations by our moving-aggregation strategy. We call our updating strategy as ”OneBP” for brevity. Fig. 2 compares our OneBP with the two backpropagation. The core idea is to cut off the gradient flow towards users’ encodings.
We detail the execution of our OneBP strategy for one user in one training epoch as follows: Let $(u_{i},v_{j},\{v_{k}\})$ denote the selected user $u_{i}$ with his/her interacted item $v_{j}$ as positive and his/her uninteracted items $\{v_{k}\}$ as negatives. Our OneBP updates $\mathbf{u}_{i}$, $\mathbf{v}_{j}$, and $\mathbf{v}_{k}$ as follows:
@@
Compared with the two backpropagation, our OneBP does not use gradient backpropagation for updating users’ representations. Instead, Eq. (6) indicates that the user encoding is updated by the aggregation over his/her previous one and his/her just-updated interacted item’s encoding in a training epoch.
@@ 4.3.2. Decoupling the Impact of False Negative Samples
We note that although contrastive learning is a powerful technique in many applications, it faces the challenge of false negative sample in the OCCF problem. That is, an unlabeled item is a user interested one, but it is not interacted simply due to not yet exposed to this user. If such an unlabeled item is selected as a negative sample for contrastive loss computation, it can lead to update user encoding in an incorrect direction. This issue has been extensively studied in the literature [38; 39; 40; 41]. Although many algorithms have been proposed to alleviate the adverse impact of false negative samples [22; 37; 7; 42; 43; 40], the existing two backpropagation cannot fully address this issue as its loss gradient is still backpropagated to update user encodings. In contrast, our OneBP only applies his/her interacted items’ encodings to update a user encoding, which can effectively decouple the impact of false negative samples.
@@ 5.1. Experiment Settings
Dataset: We conduct experiments on four widely used public datasets, including the MovieLens-100k, MovieLens-1M  [^1] MovieLens: https://grouplens.org/datasets/movielens/, Gowalla  [^2] Gowalla: https://paperswithcode.com/dataset/gowalla and Yelp2018  [^3] Yelp2018: https://www.yelp.com/dataset. Table 1 presents the statistics of the four datasets. We note that the Gowalla and Yelp2018 are often regarded as large datasets with serious interaction sparsity. In our research, we adopt a random selection process to create the training and test sets for each dataset. Specifically, we randomly select 80% of the historical interactions for each user as the training set, while the remaining interactions are treated as the test set.
@@
Competitors: We compare our proposed OneBP with the following three groups of competitors, as introduced in Section 3. (1) The first encoder group focuses on the design of user and item encoder, including the MF [32], NGCF [26], and LightGCN[8]. (2) The second negative group focuses on negative sampling yet with the simple MF encoder, including the PNS [33; 27; 34], DNS [29], HarS [5], and 2stS [6]. (3) The third loss group focuses on the loss function, including the CML [5], SFCML [13], DPCML [12], DCL [22], HCL [37], and BUIR [14].
Evaluation metrics: We adopt the mostly used performance metrics in the OCCF problem for personalized top-$K$ recommendation, including Precision (P), Recall (R), F1-Score (F1), and Normalized Discounted Cumulative Gain (NDCG). For their wide usage, we do not detail their calculations here. We use them to evaluate the top-$5$, top-$10$, and top-$20$ recommendation performance.
Parameter Settings: To make fair comparisons, we use the same number of negative samples for all algorithms. More detailed hyperparameter settings are documented in the README file of the code. The codes is implemented with Pytorch  [^4] The experimental codes of OneBP model in the paper are released at:  https://anonymous.4open.science/r/OneBP-7C87.
@@ 5.4. Hyper-parameter Study
Our OneBP contains only one hyperparameter $\beta$ to balance the weight between a user encoding in last iteration and his/her interacted item encoding updated in this iteration, c.f., Eq. (6). Note that the number of negatives $N_{s}$ is a common hyperparameter in contrastive learning. We conduct experiments to analyze the effects of the two hyperparameters.
@@ 5.5. Exploration Study
Immediately following this, we explore how well OneBP captures user interest. However, data often lack labels in the OCCF problem, making it impossible to measure the quality of learned encodings as well as the diversity of user preferences. For items without labels, we resort to the KMeans clustering algorithm [44] to cluster items’ encodings into six clusters, where one item belongs to one cluster. We use the t-SNE dimensionality reduction [45] to map items’ encodings into a two-dimensional plane, where colors indicate cluster indexes, i.e., latent types.
Fig. 4 plots the distribution of projected items’ encodings for both the training and testing dataset, where a user encoding is also projected and his/her interacted items in the training dataset are marked as rectangles and with dark green color. Fig. 4 (a) presents the distribution of items’ encodings at the random initialization phase. We can observe that items with different latent types are almost mixed and there are no clear cluster divisions. Fig. 4 (b)(c)(d) present the distribution of items’ encodings after the training of a two tower model. It can be clearly seen that items of the same color (indicating of the same latent type) are closer to each others as a cluster, and the divisions of different clusters are clearly observable. This validates our arguments to some extent, that is, a two tower recommendation model is capable of learning items’ intrinsic yet unexposed attributes as encoded in the learned items’ encodings, and items can be clustered into laten types with observable cluster divisions.
Fig. 4 (b) displays the clustering results by using one backpropagation only for the user encoding update. It can be observed that the learned items’ encodings do not present a kind of uniform dispersion in the space, and clusters’ divisions are also not that clear, as compared with Fig. 4 (c) and (d). Furthermore, the interacted items by a user in the training dataset are too much concentrated in one cluster, and the user encoding is farther away from his/her interacted items, also compared with Fig. 4 (c) and (d). These observations indicate that this user backpropagation training strategy may not be capable of learning high quality representations for items and for users, so its recommendation performance suffers.